If f(t)= eiwt at the interval (-π, π) determine fourier series of the f(t) ?


1
Expert's answer
2021-05-24T16:06:17-0400

Here the half-period is "L=\pi." Therefore, the coefficient "c_0" is


"c_0=\dfrac{1}{2\pi}\displaystyle\int_{-\pi}^{\pi}f(t)dt=\dfrac{1}{2\pi}\displaystyle\int_{-\pi}^{\pi}e^{iwt}dt"

"=\dfrac{i}{2iw\pi}(e^{iw\pi}-e^{-iw\pi})=\dfrac{\sin{(w\pi)}}{w\pi}"

For "n\not=0"


"c_n=\dfrac{1}{2\pi}\displaystyle\int_{-\pi}^{\pi}f(t)e^{-{in\pi t \over \pi}}dt=\dfrac{1}{2\pi}\displaystyle\int_{-\pi}^{\pi}e^{iwt}e^{-int}dt"

"=\dfrac{1}{2\pi i(w-n)}(e^{i(w-n)\pi}-e^{-e^{i(w-n)\pi}})"

"=\dfrac{\sin((w-n)\pi)}{ (w-n)\pi}"


"f(t)=e^{iwt}=\dfrac{\sin{(w\pi)}}{w\pi}+\displaystyle\sum_{-\infin}^{\infin}\dfrac{\sin((w-n)\pi)}{ (w-n)\pi}e^{int}"